### v1

# Examining Different Research Communities: Authorship Network thanks: My advisor Dr. Bojan Cukic

Abstract
Google Scholar is one of the top search engines to access research articles across multiple disciplines for scholarly literature. Google scholar advance search option gives the privilege to extract articles based on phrases, publishers name, authors name, time duration etc. In this work, we collected Google Scholar data (2000-2021) for two different research domains in computer science:Data Mining and Software Engineering. The scholar database resources are powerful for network analysis, data mining, and identify links between authors via authorship network. We examined coauthor-ship network for each domain and studied their network structure. Extensive experiments are performed to analyze publications trend and identifying influential authors and affiliated organizations for each domain. The network analysis shows that the networks features are distinct from one another and exhibit small communities within the influential authors of a particular domain.
Keywords:
network analysis graph theory co-author network research collaboration
1 Introduction
Social network analysis has become a popular technique in many fields such as social media networks, web networks, biological networks, academic social networks etc. The network analysis gives a broader picture of connections that has been maintained within a community over a long period of time. As academic papers are increasing exponentially, each academic area is segmented and specialized. To better understand the scientific community in general, it is important to study and model the factors of research communities, the growth and evaluation of a particular research domain. Therefore, modeling these factors efficiently is important to better understand the working culture of the scientific community. In that case, authorship network is very useful for analysis across different scientific domains. We are interested to look at the authorship network of research domains under computer science stream in recent era. The pattern and culture are quite a bit different in different research domains under same stream. Most of the authorship networks analysis finds communities and structures of sub-communities.
Many co-authorship networks have been studied to investigate collaboration among users [2, 4, 13]. Most of them used databases, for example Aminer, DBLP digital libraries to generate network. In these cases, it is difficult to separate publications for particular conferences under same stream. Co-authorship network can describe publication rate and scientific productivity [1, 11]. Previous studies have shown that, there is a strong positive relation between co-authorship network and scientific productivity [8, 5]. Researchers from diverse geographical locations [6, 12] and multi-disciplinary backgrounds collaborate together which positively impacts scientific productivity. Social network studies have furthered the understanding of the relationship between co-authorship and productivity. Studies assessing the relationship between productivity and the position of authors in the co-author network have found that authors who publish with many different co-authors bridge communication and exhibits higher rates of publication records.
In this paper, I use data collected from Google Scholar on data mining and software engineering research domain. Using this dataset, our proposed work will follow the below research contributions:
  • Investigate pattern of publication over past 20 years for different research domains.
  • Identify the most influential authors based on published articles.
  • Identify the most frequently affiliated organization appeared within a domain.
  • Analyze and compare network features of two distinct research areas.
2 Related Work
Co-authorship networks have been widely used in the past to explore scientific collaboration among authors. An authorship network is an academic social network where individuals are connected based on their collaboration through research articles. In previous studies, co-authorship networks have been analyzed to explore different statistical characteristics. It has been extensively analyzed at network level and individual node level[14] and behavioral patterns in scientific collaborations [10]. Similar to authorship network, citation network also have been studied much. In citation network, nodes connection are created from paper citations, where papers are nodes and links generate when one paper is cited by another paper. Usually citation network shows the relation among papers. On the other hand, co-authorship network shows the relation among authors. It shows a stronger bond among authors through collaboration. Co-authorship network evolves over time within research domain. Researchers from all over the world share their research interest in a specific domain or may belong to different domains. In that way, one researcher may belong to multiple co-authorship networks. Many research dimensions like finding rising stars [9], name disambiguation [7], domain experts [3] have been studied in academic social network.
In this paper, we analyze and compare the publication pattern, features of domain experts and collaboration culture of two research domains in computer science. The number of published papers per year shows how these two fields have emerged over time. From the google scholar profile, it is easy to capture authors information and affiliations to understand the most involved organizations/institutions in a domain. Also, the network analysis shows relation among authors and degree distribution of authors.
3 Data Collection
The dataset used in this work is collected from Google Scholar. The data is collected for two different research domains: Data Mining and Software Engineering. For each research domain, three recognized and popular conferences have been chosen to extract using advanced search option. The three selected conferences for Data Mining are, 1. IEEE International Conference on Machine Learning and Applications (ICMLA) 2. IEEE International Conference on Data Mining (ICDM) 3. ACM International Conference on Knowledge Discovery and Data Mining (SIGKDD). On the other hand, for Software Engineering the selected conferences are: 1. International Conference on Software Engineering (ICSE) 2. ACM International Symposium on Foundations of Software Engineering (SIGSOFT) 3. IEEE/ACM International Conference on Automated Software Engineering (ASE). For each conference, we have extracted 1000 published papers within time-frame 2000 to 2021. In total, we have collected 3000 published papers for each area.
A published paper has one or more than one authors involved in it. We have listed authors’ information from the papers. We can identify 2482 authors in Software Engineering conferences and 3787 authors in Data Mining conferences. We have collected the author’s information: full name, citation, h-index, and affiliation from their Google Scholar profile. We have used this information for further analysis.
4 Methods and Results
The research article summarizes the idea of the article. It describes the purpose and content of the research paper adequately in fewer words and definitely the first read. To identify the most frequently used words in the title of the articles of a domain, we have created a word cloud based on frequently used listed words. Figure 1 shows the word clouds of data mining and software engineering based on research article titles. Following these words, data mining research communities seem interested in using deep learning approaches for prediction-based tasks for example classification. On the other hand, software engineering communities are more focused on systems, security, and testing for analysis.
Figure 1: Wordcloud of Data Mining (left) and Software Engineering (right)
4.1 Paper Publication Pattern
In order to understand the publication record of these two domains, we have accumulated the number of published papers each year within the time frame. The trend shows how the two domains are different in terms of paper publishing each year. Figure 2 shows that data mining publication records have increased over the years which means researchers have paid more attention to this domain. On the other hand, the record has changed and reduced over time for software engineering. This information definitely shows how one research area is emerging more than the other. From the distribution, we can see that for data mining conferences, publication significantly increased from 2012 and the highest number of published papers was 312 in the year 2018. On the other hand, software engineering reached its highest peak in 2005 and after that, the number of published papers/acceptance rates slowly decreased every year. The software engineering conferences had highest number of publications 238 in 2005 and the lowest in the year 2020 and 2021.
Figure 2: Number of publications per year: The left image shows Data Mining and right image shows Software Engineering
4.2 Influential Authors
Authors publish papers at the same conferences every year. From the number of total published papers within the research domain, we can mark the top 10 influential authors. In our consideration, the author who published papers the highest time as first author or as co-author within a domain is the most influential one. We have identified the top 10 influential authors based on their publication records in our dataset.
From figure 3, we can see the top 5 influential authors in data mining are Jiawei Han, Huan Liu, Eamonn Keogh, Philip Yu, and Ryan Baker. On the other hand, the top 5 influential authors in software engineering domain are Barbara Kitchenham, Thomas Zimmermann, Mark Harman, Gail Murphy, and Krysztof Czarnecki. Here, Jiawei Han and Huan Liu from data mining have appeared total 32 and 30 times accordingly, whereas Barbara Kitchenham and Thomas Zimmermann from software engineering have appeared 35 and 26 times accordingly within time frame 2000 to 2021.
Figure 3: The top 10 influential authors in Data Mining
Figure 4: The top 10 influential authors in Software Engineering
4.3 Affiliated Organizations
We also looked at the authors Google Scholar profiles that are classified as valid profiles. The valid profiles usually have full name, research interest, citations, h-index, and affiliated organization information. We have captured the organizations name of the authors. From that we have come up with the top most frequently appeared affiliated organizations of the authors for each research domain. In data mining domain, we can see that Google has been the topmost affiliated company to collaborated with, then University of Minnesota, Microsoft, University of Illinois Urbana-Champaige, University of North California, Riverside so and so forth. In software engineering, Microsoft has been the topmost affiliated company, and then Keele University, Carnegie Mellon University, Facebook, Waterloo University so and so forth (Figure 5).
Figure 5: Number of publications per year: The left image shows Data Mining and right image shows Software Engineering
5 Network Analysis
In this section, we perform a co-author network analysis to examine the relationships between the authors within domains. We have generated undirected network for both domains. Data mining co-author network has 3787 nodes and 4363 edges. There are 3538 connected components in the network and the giant connected component has 982 nodes involved within it. Figure 6 shows the largest connected component on the top and degree distribution for all the nodes. The distribution shows that more than 1200 nodes have degree value of 3 and few nodes have degree value above 10. We found Huan Liu has worked with 41 authors and Jiawei Han has worked with 39 authors.
On the other hand, the software engineering co-author net has 2482 nodes and 2870 edges. Out of the whole network, there are 572 connected components. The giant connected component of the network consists of 708 nodes. From the figure 7, we can see the largest connected component of 708 nodes. The degree histogram (left bottom figure 7) of the whole network shows that more than 800 nodes have degree 2 and few nodes have degree 10. In this case, David Lo and Mark Harman had worked with 20 and 19 authors correspondingly.
Compared to software engineering, data mining authors’ highest degree value 40 is twice that of software engineering authors’ highest degree value 20.
Figure 6: Degree Distribution of Software Engineering Authors
Figure 7: Degree Distribution of Software Engineering Authors
5.1 Collaboration Teams
In today’s research system, team collaboration is the most essential part. In a research collaboration team, collaborators interchange their ideas, thoughts, datasets, problem-solving approaches, writing, etc. Some of the effective teams work together many times. It is important to point out strong tied collaboration teams in the co-author network to understand collaboration culture in a research domain.
From the network analysis, we have brought out the paired collaborators who co-authored together several times. We found that in software engineering, Barbara Kitchenham has collaborated with Pearl Brereton highest 13 times and with David Budgen 12 times. Also, ”Per Runeson and Martin Höst”, ”Nachiappan Nagappan and Thomas Ball”, ”Andrea De Lucia and Rocco Oliveto” each paired collaborators worked together 8 times within the time-frame for the conferences mentioned early.
On the other hand, in data mining ”Cristóbal Romero and Sebastián Ventura” collaborated together highest 12 times, ”Jiliang Tang and Huan Liu” 7 times. ”George Karypis and Joseph Konstan”, ”Eamonn Keogh and Stefano Lonardi”, ”Igor Santos and Xabier Ugarte-Pedrero” each paired collaborator worked together 5 times.
5.2 Connected Authors
From a high level, the connected components have several key features. Connected components can be interpreted as small communities generated in the network. We focus on the second and third largest connected clusters of the network. To ease our understanding, we have labeled nodes with green color when authors have citations more than 1000 times, blue when above 500 times, authors that we could not capture the citation number are labeled as yellow and rest of the nodes are labeled as red. The node sizes are defined by the number of times authors are cited in total from their Google Scholar profiles. And edge sizes are defined by the number of times two authors have been collaborated.
5.2.1 Findings
The three clusters from two domains have distinct characteristics. Cluster A is small community with relatively less frequent collaboration among authors than Cluster B, There is also less number of beginner researcher compared to B. On the other hand, cluster C is combination of high level researcher to beginner level researchers including frequent collaboration information.
5.2.2 Cluster A: Data Mining third largest cluster
Cluster - A (figure 8 left) third largest connected cluster from data mining network consists of 12 nodes who are interconnected. It is common in scientific communities that highly cited authors work with authors who has strong publication records. From cluster A we can see that, most of the author are connected with other authors who have publications over 100 times. In this community not much beginner level researchers are involved.
Figure 8: Connected Components of Data Mining - Cluster A (left) and Cluster B (right)
5.2.3 Cluster B: Data Mining third largest cluster
Cluster - B (figure 8 right) second largest connected cluster consists of 14 authors from data mining network. Here the biggest node is author Michael Jordan from UC Berkeley, seem to be closely connected with very beginner level researchers and other domain expert authors simultaneously. Also, he has strong ties with beginner researchers and domain experts.
5.2.4 Cluster C: Software Engineering second largest cluster
Cluster - C (figure 9) the second largest connected component of software engineering consists of 21 authors who are closely connected. Here, a big group of beginner level researchers have co-author together multiple times. Also we can see that there are couple of triangles have formed who have worked together multiple times.
Figure 9: Connected Component of Software Engineering - Cluster C
6 Conclusion
We studied the characteristics of two different research communities: data mining and software engineering authorship network from google scholar. Analyzing network structure, collaboration and author profiles of these two communities allowed us to discover interesting features. We have observed that data mining research domain has flourished more than software engineering over past 10 years. Though we have seen that the influential authors have similar number of publication record in both research domain, however, in data mining have more researchers compared to software engineering. We have seen that from the network analysis that data mining author have more collaboration than software engineering authors. The overall collaboration culture and number of authors helped to flourish data mining domain. Furthermore, the giant tech company for example Google collaborators seemed to be more interested in data mining domain and also Microsoft collaborators have higher publication record in data mining than software engineering tracks.
We would like to mention here that we do not intend to generalize the findings of this work because only selective conferences data have been used. However, we do believe that it shows an overall view of the current state-of-the-art. We wanted to present findings and demonstrate it through out dataset. This process can be extended to many directions in future. Authorship network can be useful to predict future collaboration teams. We are highly interested to explore citations network of each domain and analyze citation authors and articles network. We were able to extract authors google scholar profile information, which can be useful to retain latent information about gender and racial biases for each domain.